# Long-lived oscillations of false and true vacuum states in neutral atom systems

### v1

@@ Abstract
Metastable false vacuum states arise in a range of quantum systems and can be observed in various dynamical scenarios, including decay, bubble nucleation, and long-lived oscillations. False vacuum phenomenology has been examined in quantum many-body systems, notably in 1D ferromagnetic Ising spin systems and superfluids. In this paper, we study long-lived oscillations of false and true vacuum states in 1D antiferromagnetic neutral atom chains with long-range Rydberg interactions. We use a staggered local detuning field to achieve confinement. Using theoretical and numerical models, we identify novel spectral signatures of quasiparticle oscillations distinct to antiferromagnetic neutral atom systems and interpret them using a classical energy model of deconfinement from Rydberg tails. Finally, we evaluate the experimental accessibility of our proposed setup on current neutral-atom platforms and discuss experimental feasibility and constraints.
@@ I Introduction
A false vacuum is a metastable excited state that can arise in quantum field theory and quantum many-body physics, separated in energy from the global true vacuum. These states were explored in statistical physics [1, 2], and then prominently in quantum field theory [3, 4, 5, 6, 7, 8, 9]. False and true vacua can be identified and distinguished by the signatures they produce during dynamical processes, including false vacuum decay, bubble nucleation, and long-lived oscillations. The detection of these vacuum states can serve as a probe of the structure of the underlying system and, by extension, its history.
@@
Phase transitions with metastable false vacua are also prominent processes in quantum many-body systems. In ultracold atom systems, a number of experiments have studied such phase transitions [14, 15, 16], including recent work that detected false vacuum decay using ferromagnetic superfluids formed from atomic Bose-Einstein condensates [16], and many others have been proposed (e.g. [17, 18, 19, 20, 21] and references therein). In 1D Ising spin systems, an array of analytic and numerical efforts have probed phase transitions involving false vacua, aided by an improved understanding of confinement physics in these systems [22, 23, 24, 25]. The rate of false vacuum decay has been derived and quantified in detail [26, 27]. Numerical studies have shown that a linear ramp across a first order phase transition can produce nucleation events in a false vacuum background at quantized values of the ramp field [28]. Simulations of false vacuum domain wall collisions have shown that confined bubbles can be produced in these scattering events [29].
Instead of phase transitions, false and true vacua can also be detected through long-lived oscillations arising from superpositions with excited states containing subcritical domains [30, 31]. In contrast to decay, long-lived oscillations are readily observable and do not depend on a vanishingly small occurrence probability. In 1D Ising spin chains, recent numerical work has examined the Fourier spectra of oscillations after a quantum quench, finding spectral trends that can be used to identify and discriminate between false and true vacua [31]. This followed earlier work that observed Bloch oscillations of domain walls after a quench, producing a regular peak structure in the Fourier spectra [30]. In neutral atom chains, false and true vacuum oscillations have been briefly examined in a supplementary study, under some simplifying conditions [31].
@@
We study the long-lived oscillations of false and true vacuum states in 1D neutral atom chains mediated by Rydberg interactions. Our main tools are numerical simulations and analytic modeling. We perform a broad investigation of the parameter space and include long-range Rydberg tails, distinct from the supplementary study of Ref. [31]. Our setups are accessible to near-term experiments. We determine the spectral features of quasiparticle oscillations involving false and true vacua. We identify features distinct to these systems and explain them through deconfinement effects from Rydberg interaction tails. Our results show that the detailed microphysics of the system influence the spectral features, and thus the ability to distinguish between false and true vacuum states.
In a concurrent work, we study the dynamics of false vacuum decay and nucleation in neutral atom systems [69]. The parameter regime for that process is complementary to the one examined in this paper.
The remainder of our paper is organized as follows. We describe our model in Section II. We summarize some theoretical background for modeling the dynamics of single bubbles in Section III. Our numerical setup and results are presented in Section IV. We interpret our results in terms of a deconfinement model in Section V. In Section VI, we discuss the experimental accessibility of our setups and results. We conclude and present our outlook in Section VII.
@@ II Model
Figure 1: A schematic overview of our system and results. The atoms are arranged in a circle with 16 sites and constant atom separation. The blue +/+/-+ / - signs show the spatial pattern of the staggered local detuning field $\Delta_{\mathrm{loc},j}$ that we apply from Equation 3, here shown for $\Delta_{\mathrm{loc}}>0$. The field produces an asymmetric potential with false vacuum (FV) and true vacuum (TV) states separated by an energy density $\mathcal{E}\propto\Delta_{\mathrm{loc}}$. The vacuum states are dominated by the two $Z_{2}$ product states, shown in the chains above, where a red atom indicates a Rydberg state $\left\lvert r\right\rangle$ and a white atom indicates a ground state $\left\lvert g\right\rangle$. The false vacuum undergoes coherent oscillations with low-lying excited states, only schematically shown here inside the left well. The magnetization is a suitable observable, defined in Equations (7) and (8). Its Fourier spectrum is shown in the top right panel, exhibiting a distinctive peak structure. Similarly, the true vacuum undergoes coherent oscillations, and the associated spectrum is shown in the bottom right panel. The peak structure depends on the microscopic details of the confining potential. It can be used to distinguish between the two vacuum states, and understood from our analysis in Section V.
We study long-lived oscillations of an initial state near the false and true vacuum states. Figure 1 schematically shows the dynamics and signatures in these two cases, which will be elucidated in Sections III and IV. In our numerical studies, we set the initial state near one of the vacua and track its evolution using various observables. In this section, we describe our system and observables.
@@
then we split the ground state into false and true vacua separated by energy density $\mathcal{E}\propto\Delta_{\mathrm{loc}}$ (Figure 1). We can specify the detuning using two dimensionless free parameters
@@
where we take $\alpha>0$. The $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$ will be near the false vacuum for small $\beta>0$ and near the true vacuum for small $\beta<0$. Figure 1 schematically shows the dynamics and signatures for an initial state near each vacua, though we note that the approach of the figure is the reverse of the approach we just described, in that the figure shows only $\beta>0$ and instead selects different initial states to achieve proximity to the false and true vacua.
The staggered field introduces a confining potential between domain walls that scales linearly with the domain size. The antiferromagnetic domain walls correspond to adjacent $00$ or $11$ configurations, though $11$ domain walls are energetically suppressed when the nearest-neighbor Rydberg blockade is large. The domain, or bubble, size is given by the distance between its domain walls. Domains of even size are suppressed due to the nearest-neighbor Rydberg blockade. The dynamics of domain walls can be effectively described in terms of the dynamics of quasiparticles, using concepts borrowed from high-energy physics. Domain walls can be understood as quarks and antiquarks, depending on the transition direction. In the staggered field, false vacuum domains experience a confining interaction potential and thus correspond to mesons, i.e. bound two-quark quasiparticles; true vacuum domains experience an anti-confining interaction potential. A state with domains contains at least one meson state, possibly more.
@@
which simply subtracts the mean $\bar{M}$. Locally, an evolved state has domain walls separating true and false vacua. The density of domain walls $\rho=\langle\hat{R}\rangle$ is determined from the operator
@@ III Quasiparticle Oscillations
False and true vacua can be detected and characterized using an analogous procedure based on these insights. In particular, a quench tuned to the false or true vacuum will produce a post-quench state that oscillates in superposition between that vacuum state and proximate low-energy single meson states. The post-quench oscillations will yield a Fourier spectrum that encodes the vacuum properties through the interference with the meson states. The dynamics are more easily isolated when supercritical bubbles are energetically suppressed and the mesons are subcritical.
In our numerical studies in Section IV, we initialize the state to one of the vacuum states, evolve it under the Hamiltonian, and examine the Fourier spectra of the magnetization oscillations. In this section, we present a perturbative model for examining quasiparticle excitations that provides an interpretive framework for our numerical results. We first review earlier work in spin systems. We then present our efforts with neutral atoms.
@@ III.1 Spin Chains
False and true vacuum oscillations have previously been examined using quench dynamics in the ferromagnetic 1D Ising spin model. Notably, in an extensive study, Lagnese et al. [31] performed quenches to initial states near either the false or true vacuum, and examined the Fourier spectrum of the magnetization for the resulting long-lived oscillations. They observed several features: the spectra exhibit a cluster of peaks; for an initial state near the false vacuum, the peaks increase in amplitude and spacing with increasing frequency; for an initial state near the true vacuum, the peaks decrease in amplitude and spacing with increasing frequency; and in both cases, the peaks shift to lower frequencies and spread out with increasing confinement parameter. In earlier work, Pomponio et al. [30] studied a different type of false vacuum oscillation. They performed an anti-confining quench and unexpectedly observed long-lived oscillations, interpreted as Bloch oscillations of the domain walls. The Fourier spectra of the correlation functions exhibited a regular peak structure.
Lagnese et al. [31] accurately calculated the frequencies and amplitudes of the spectral peaks by restricting to the manifold of low-energy 0- and 1-meson states and examining the dynamics using perturbation theory. This followed Rutkevich [26] who used a similar approach to compute the tunneling probability in false vacuum decay. We briefly summarize the model developed in these studies. It can be used more generally for any condensed matter system with confinement, and we thus use a more general notation in parts. In Section III.2, we modify it to examine the dynamics in our neutral atom system.
@@
The perturbation conserves momentum $k$; at first order, it thus connects the vacuum states, which have $k=0$, to the 1-meson eigenstates with $k=0$, and only at higher orders to even superpositions of 1-meson eigenstates with degenerate energies and opposite momentum. To compute the 1-meson eigenstates with $k=0$, one can project the Schrödinger equation for $\tilde{H}_{0}$ to the 1-meson sector, take the coordinate representation, and shift to the center of momentum frame $k=0$, obtaining [26, 31]
@@ III.2 Neutral Atom Setup
The evolved state $\left\lvert\psi(t)\right\rangle$ can be approximated using standard techniques from time-dependent perturbation theory. The initial state $\left\lvert\psi(0)\right\rangle$ is approximately the desired vacuum state of $\tilde{H}_{0}$, which can be denoted $\left\lvert 0_{v}\right\rangle$ with eigenvalue $\varepsilon_{v}$ where $v=f,t$ labels the vacuum type. To first order in perturbation theory, as examined in Ref. [31], the evolved state is
@@
where $d_{\ell}^{*}(t)=(1-e^{-i\tilde{\omega}_{\ell v}t})/(i\tilde{\omega}_{\ell v})$. The major peaks in the Fourier spectrum of $M$ can be understood from this expression. The frequency of a major peak equals a frequency $\tilde{\omega}_{\ell v}=\varepsilon_{\ell}-\varepsilon_{v}$ between one of the meson states $\left\lvert\phi_{\ell}\right\rangle$ and the vacuum state $\left\lvert 0_{v}\right\rangle$. The energies $\varepsilon_{v}$ and $\varepsilon_{\ell}$ are the eigenvalues of the unperturbed Hamiltonian $\tilde{H}_{0}$; however, since the perturbation is weak, these are approximately equal to the eigenvalues of the exact Hamiltonian $H$. The amplitude of the peak depends on the two matrix elements $\langle\phi_{\ell}|\tilde{U}|0_{v}\rangle$ and $\langle 0_{v}|\hat{M}|\phi_{\ell}\rangle$. The Fourier spectrum of $M$ can thus probe interference between the desired vacuum state and proximate 1-meson states, connected by $\tilde{U}$ and $\hat{M}$, which can potentially identify the vacuum.
In rough terms, the Hamming distance of the state $\left\lvert\phi_{\ell}\right\rangle$ from the false vacuum state $\left\lvert 0_{v}\right\rangle$, dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, is inversely correlated with the size of the matrix elements $\langle\phi_{\ell}|\tilde{U}|0_{v}\rangle$ and $\langle 0_{v}|\hat{M}|\phi_{\ell}\rangle$. This can be explained as follows. The energies of the 1-meson eigenstates are inversely correlated with the domain size for a false vacuum initial state, and a subset are correlated for a true vacuum initial state (Section V examines these trends in detail). As the domain size increases, the non-zero terms in the matrix elements arise from higher order domain wall hoppings. In lieu of directly computing the matrix elements, we use the Hamming distance to estimate trends.
@@ IV Numerical Study
Figure 2: The numerical setup to study false and true vacuum long-lived oscillations. The left section shows the atom configuration. The grey circle shows the blockade radius. The initial state is set to the ground state of $H_{\mathrm{prep}}=H(\beta_{\mathrm{prep}})$ with $\beta_{\mathrm{prep}}=-10^{-3}$; it is dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, represented by the color pattern for the atoms. The state then evolves under $H$ with a different $\beta$. The right section shows the waveform sequence. The waveforms are all held constant. The local detuning field for a small positive value $\beta\propto\Delta_{\mathrm{loc}}>0$ is shown by the blue outer +/+/-+ / - symbols on the left and the labeled waveform on the right, producing an initial state close to the false vacuum state of $H$. The field for a small negative value $\beta\propto\Delta_{\mathrm{loc}}<0$ is shown by the teal inner +/+/-+ / - symbols on the left and the labeled waveform on the right, producing an initial state close to the true vacuum state of $H$.
We study the long-lived oscillations of false and true vacua in 1D neutral atom systems, using the setup shown in Figure 2. The atoms lie in a ring with $n_{s}$ sites, where we choose $n_{s}=16$. In particular, we use a staggered local detuning field to achieve confinement, yielding a confinement energy density $\mathcal{E}\propto\Delta_{\mathrm{loc}}$. We set the initial state $\left\lvert\psi(0)\right\rangle$ to be the ground state of $H_{\mathrm{prep}}=H(\beta_{\mathrm{prep}})$, obtained from exact diagonalization, where $H$ is the Rydberg Hamiltonian in Equation 1 and $\beta_{\mathrm{prep}}=-10^{-3}$. It is dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, represented by the colors in the atom chain in Figure 2. We evolve the state under $H$ with a different local detuning parameter $\beta$, which has a small constant amplitude $|\beta|\ll 1$. In effect, we perform a quench $\delta\beta=\beta-\beta_{\mathrm{prep}}$. The initial state is then close to the false vacuum of $H$ for $\beta\propto\Delta_{\mathrm{loc}}>0$, shown by the blue outer +/+/-+ / - pattern and waveform in Figure 2; it is close to the true vacuum of $H$ for $\beta\propto\Delta_{\mathrm{loc}}<0$, shown by the teal inner +/+/-+ / - pattern and waveform. We note that in our procedure here, we hold the initial state fixed and modify the spatial pattern to obtain proximity to the false and true vacua; this is the reverse of the schematic in Figure 1, where the spatial pattern is held fixed and the different states show the proximity.
@@
Figure 3: The oscillation properties of the magnetization $M$ for $\beta=\pm 0.01$. (a) The magnetization $M$ vs (dimensionless) time $\Omega t/2\pi$ for $\beta=0.01$. (b) The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ of the transformed magnetization $M_{T}$, defined in Equation 8, for $\beta=0.01$, which yields an initial state close to the false vacuum. The curves with different opacities show results for $M_{T}$ evolved to different stop times $t_{\mathrm{Stop}}$. (c) The simultaneous eigenstates of $H$ and $T_{2}$. The eigenstates are identical for $\beta=\pm 0.01$, by symmetry. The true vacuum energy is set to zero, $E_{t}\rightarrow 0$, as indicated in the bottom left. The color bar shows the domain wall density $\rho n_{s}$; the blue points show the false vacuum (FV) and true vacuum (TV) eigenstates, and the red points show the $1$-meson eigenstates. (d) The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ for $\beta=-0.01$, which yields an initial state close to the true vacuum. In (b) and (d), the vertical grey lines show the frequencies $\omega_{\ell v}=E_{\ell}-E_{v}$, where $E_{\ell}$ for $\ell=1,\ldots,8$ are the energy eigenvalues of the 1-meson eigenstates with momentum $k=0$ and $E_{v}$ is the energy eigenvalue of the appropriate vacuum state $v=f,t$. The Fourier spectra for $\beta=\pm 0.01$ both have a similar dominant peak but different clusters of secondary peaks, which can be used to distinguish between the false and true vacua, and understood from our analysis in Section V.

### v2

# Long-lived oscillations of metastable states in neutral atom systems

@@ Abstract
Metastable states arise in a range of quantum systems and can be observed in various dynamical scenarios, including decay, bubble nucleation, and long-lived oscillations. The phenomenology of metastable states has been examined in quantum many-body systems, notably in 1D ferromagnetic Ising spin systems and superfluids. In this paper, we study long-lived oscillations of metastable and ground states in 1D antiferromagnetic neutral atom chains with long-range Rydberg interactions. We use a staggered local detuning field to achieve confinement. Using theoretical and numerical models, we identify novel spectral signatures of quasiparticle oscillations distinct to antiferromagnetic neutral atom systems and interpret them using a classical energy model of short-range meson repulsion. Finally, we evaluate the experimental accessibility of our proposed setup on current neutral-atom platforms and discuss experimental feasibility and constraints.
@@ I Introduction
Metastable states are excited states that can arise in quantum field theory and quantum many-body physics, separated in energy from the global ground state. These states were explored in statistical physics [1, 2], and then prominently in quantum field theory in the form of false vacua [3, 4, 5, 6, 7, 8, 9]. Metastable and ground states can be identified and distinguished by the signatures they produce during dynamical processes, including decay, bubble nucleation, and long-lived oscillations. The detection of these states can serve as a probe of the structure of the underlying system and, by extension, its history.
@@
Phase transitions with metastable states are also prominent processes in quantum many-body systems. In ultracold atomic systems, a number of experiments have studied such phase transitions [14, 15, 16], including recent work that detected metastable decay using ferromagnetic superfluids formed from atomic Bose-Einstein condensates [16], and many others have been proposed (e.g. [17, 18, 19, 20, 21] and references therein). In 1D Ising spin systems, an array of analytic and numerical efforts have probed phase transitions involving metastability, aided by an improved understanding of confinement physics in these systems [22, 23, 24, 25]. The rate of decay has been derived and quantified in detail [26, 27]. Numerical studies have shown that a linear ramp across a first order phase transition can produce nucleation events in the background state at quantized values of the ramp field [28]. Simulations of domain wall collisions have shown that confined domains can be produced in these scattering events [29].
Instead of phase transitions, metastable and ground states can also be detected through long-lived oscillations arising from superpositions with excited states containing subcritical domains [30, 31]. In contrast to decay, long-lived oscillations are readily observable and do not depend on a vanishingly small occurrence probability. In 1D Ising spin chains, recent numerical work has examined the Fourier spectra of oscillations after a quantum quench, finding spectral trends that can be used to identify and discriminate between metastable and ground states [31]. This followed earlier work that observed Bloch oscillations of domain walls after a quench, producing a regular peak structure in the Fourier spectra [30]. In neutral atom chains, oscillations have been briefly examined in a supplementary study, under some simplifying conditions [31]. Long-lived oscillations can also arise in quantum many-body systems under conditions unrelated to metastability, and can be similarly used to interrogate the system properties (e.g. [32] and references therein).
False and true vacuum phenomenology in quantum field theory is conceptually similar to metastable and ground state phenomenology in truncated quantum many-body systems, but differs fundamentally since the vacuum state of a quantum field includes all field modes whereas the states treated in a finite-dimensional Hilbert space are comparatively restricted. Metastable states have intrinsically interesting dynamics in condensed matter systems, and though observations on these systems cannot be directly applied to quantum fields, they can interrogate the emergence of new phenomena in truncated systems and potentially inspire further theoretical studies.
@@
We study the long-lived oscillations of metastable and ground states in 1D neutral atom chains mediated by Rydberg interactions. Our main tools are numerical simulations and analytic modeling. We perform a broad investigation of the parameter space and include long-range Rydberg tails, distinct from the supplementary study of Ref. [31]. Our setups are accessible to near-term experiments. We determine the spectral features of quasiparticle oscillations involving metastable and ground states. We identify features distinct to these systems and explain them through the competition between confinement and short-range repulsive effects. Our results show that the detailed microphysics of the system influence the spectral features, and thus the ability to distinguish between metastable and ground states.
In a concurrent work, we study the decay and nucleation dynamics of metastable excited states in neutral atom systems [70]. The parameter regime for that process is complementary to the one examined in this paper.
The remainder of our paper is organized as follows. We describe our model in Section II. We summarize some theoretical background for modeling the dynamics of single bubbles in Section III. Our numerical setup and results are presented in Section IV. We interpret our results in terms of a short-range meson repulsion model in Section V. In Section VI, we discuss the experimental accessibility of our setups and results. We conclude and present our outlook in Section VII.
@@ II Model
Figure 1: A schematic overview of our system and results. The atoms are arranged in a circle with 16 sites and constant atom separation. The blue +/+/-+ / - signs show the spatial pattern of the staggered local detuning field $\Delta_{\mathrm{loc},j}$ that we apply from Equation 3, here shown for $\Delta_{\mathrm{loc}}>0$. The field produces an asymmetric potential with metastable state (MS) and ground state (GS) separated by an energy density $\mathcal{E}\propto\Delta_{\mathrm{loc}}$. The states are dominated by the two $Z_{2}$ product states, shown in the chains above, where a red atom indicates a Rydberg state $\left\lvert r\right\rangle$ and a white atom indicates a ground state $\left\lvert g\right\rangle$. The metastable state undergoes coherent oscillations with low-lying excited states, only schematically shown here inside the left well. The magnetization is a suitable observable, defined in Equations (7) and (8). Its Fourier spectrum is shown in the top right panel, exhibiting a distinctive peak structure. Similarly, the ground state undergoes coherent oscillations, and the associated spectrum is shown in the bottom right panel. The peak structure depends on the microscopic details of the confining potential. It can be used to distinguish between the two states, and understood from our analysis in Section V.
We study long-lived oscillations of an initial state near the metastable and ground states. Figure 1 schematically shows the dynamics and signatures in these two cases, which will be elucidated in Sections III and IV. In our numerical studies, we set the initial state near one of the states and track its evolution using various observables. In this section, we describe our system and observables.
@@
then we split the ground state into a metastable and ground state separated by energy density $\mathcal{E}\propto\Delta_{\mathrm{loc}}$ (Figure 1). We can specify the detuning using two dimensionless free parameters
@@
where we take $\alpha>0$. The $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$ will be near the metastable state for small $\beta>0$ and near the ground state for small $\beta<0$. Figure 1 schematically shows the dynamics and signatures for an initial state in each case, though we note that the approach of the figure is the reverse of the approach we just described, in that the figure shows only $\beta>0$ and instead selects different initial states to achieve proximity to the metastable and ground states.
The staggered field introduces a confining potential between domain walls that scales linearly with the domain size. The antiferromagnetic domain walls correspond to adjacent $00$ or $11$ configurations, though $11$ domain walls are energetically suppressed when the nearest-neighbor Rydberg blockade is large. The domain, or bubble, size is given by the distance between its domain walls. Domains of even size are suppressed due to the nearest-neighbor Rydberg blockade. The dynamics of domain walls can be effectively described in terms of the dynamics of quasiparticles, using concepts borrowed from high-energy physics. Domain walls can be understood as quarks and antiquarks, depending on the transition direction. In the staggered field, metastable state domains experience a confining interaction potential and thus correspond to mesons, i.e. bound two-quark quasiparticles; ground state domains experience an anti-confining interaction potential. A state with domains contains at least one meson state, possibly more.
@@
which simply subtracts the mean $\bar{M}$. Locally, an evolved state has domain walls separating ground and metastable states. The density of domain walls $\rho=\langle\hat{R}\rangle$ is determined from the operator
@@ III Quasiparticle Oscillations
Metastable and ground states can be detected and characterized using an analogous procedure based on these insights. In particular, a quench tuned to either state will produce a superposition between that state and proximate low-energy single meson states. The post-quench oscillations will yield a Fourier spectrum that encodes the metastable or ground state properties through the interference with the meson states. The dynamics are more easily isolated when supercritical bubbles are energetically suppressed and the mesons are subcritical.
In our numerical studies in Section IV, we initialize the state near the metastable or ground state, evolve it under the Hamiltonian, and examine the Fourier spectra of the magnetization oscillations. In this section, we present a perturbative model for examining quasiparticle excitations that provides an interpretive framework for our numerical results. We first review earlier work in spin systems. We then present our efforts with neutral atoms.
@@ III.1 Spin Chains
Metastable oscillations have previously been examined using quench dynamics in the ferromagnetic 1D Ising spin model. Notably, in an extensive study, Lagnese et al. [31] performed quenches to initial states near either the metastable or ground state, and examined the Fourier spectrum of the magnetization for the resulting long-lived oscillations. They observed several features: the spectra exhibit a cluster of peaks; for an initial state near the metastable state, the peaks increase in amplitude and spacing with increasing frequency; for an initial state near the ground state, the peaks decrease in amplitude and spacing with increasing frequency; and in both cases, the peaks shift to lower frequencies and spread out with increasing confinement parameter. In earlier work, Pomponio et al. [30] studied a different type of metastable oscillation. They performed an anti-confining quench and unexpectedly observed long-lived oscillations, interpreted as Bloch oscillations of the domain walls. The Fourier spectra of the correlation functions exhibited a regular peak structure.
Lagnese et al. [31] accurately calculated the frequencies and amplitudes of the spectral peaks by restricting to the manifold of low-energy 0- and 1-meson states and examining the dynamics using perturbation theory. This followed Rutkevich [26] who used a similar approach to compute the tunneling probability in the decay of metastable excited states. We briefly summarize the model developed in these studies. It can be used more generally for any condensed matter system with confinement, and we thus use a more general notation in parts. In Section III.2, we modify it to examine the dynamics in our neutral atom system.
@@
The perturbation conserves momentum $k$; at first order, it thus connects the metastable and ground states, which have $k=0$, to the 1-meson eigenstates with $k=0$, and only at higher orders to even superpositions of 1-meson eigenstates with degenerate energies and opposite momentum. To compute the 1-meson eigenstates with $k=0$, one can project the Schrödinger equation for $\tilde{H}_{0}$ to the 1-meson sector, take the coordinate representation, and shift to the center of momentum frame $k=0$, obtaining [26, 31]
@@ III.2 Neutral Atom Setup
The evolved state $\left\lvert\psi(t)\right\rangle$ can be approximated using standard techniques from time-dependent perturbation theory. The initial state $\left\lvert\psi(0)\right\rangle$ is approximately the desired metastable or ground state of $\tilde{H}_{0}$, which can be denoted $\left\lvert 0_{v}\right\rangle$ with eigenvalue $\varepsilon_{v}$ where $v=m,g$ labels the state type. To first order in perturbation theory, as examined in Ref. [31], the evolved state is
@@
where $d_{\ell}^{*}(t)=(1-e^{-i\tilde{\omega}_{\ell v}t})/(i\tilde{\omega}_{\ell v})$. The major peaks in the Fourier spectrum of $M$ can be understood from this expression. The frequency of a major peak equals a frequency $\tilde{\omega}_{\ell v}=\varepsilon_{\ell}-\varepsilon_{v}$ between one of the meson states $\left\lvert\phi_{\ell}\right\rangle$ and the state $\left\lvert 0_{v}\right\rangle$. The energies $\varepsilon_{v}$ and $\varepsilon_{\ell}$ are the eigenvalues of the unperturbed Hamiltonian $\tilde{H}_{0}$; however, since the perturbation is weak, these are approximately equal to the eigenvalues of the exact Hamiltonian $H$. The amplitude of the peak depends on the two matrix elements $\langle\phi_{\ell}|\tilde{U}|0_{v}\rangle$ and $\langle 0_{v}|\hat{M}|\phi_{\ell}\rangle$. The Fourier spectrum of $M$ can thus probe interference between the desired state and proximate 1-meson states, connected by $\tilde{U}$ and $\hat{M}$, which can potentially identify it.
In rough terms, the Hamming distance of the state $\left\lvert\phi_{\ell}\right\rangle$ from the state $\left\lvert 0_{v}\right\rangle$, dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, is inversely correlated with the size of the matrix elements $\langle\phi_{\ell}|\tilde{U}|0_{v}\rangle$ and $\langle 0_{v}|\hat{M}|\phi_{\ell}\rangle$. This can be explained as follows. The energies of the 1-meson eigenstates are inversely correlated with the domain size for a metastable initial state, and a subset are correlated for a ground initial state (Section V examines these trends in detail). As the domain size increases, the non-zero terms in the matrix elements arise from higher order domain wall hoppings. In lieu of directly computing the matrix elements, we use the Hamming distance to estimate trends.
@@ IV Numerical Study
Figure 2: The numerical setup to study metastable and ground state long-lived oscillations. The left section shows the atom configuration. The grey circle shows the blockade radius. The initial state is set to the ground state of $H_{\mathrm{prep}}=H(\beta_{\mathrm{prep}})$ with $\beta_{\mathrm{prep}}=-10^{-3}$; it is dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, represented by the color pattern for the atoms. The state then evolves under $H$ with a different $\beta$. The right section shows the waveform sequence. The waveforms are all held constant. The local detuning field for a small positive value $\beta\propto\Delta_{\mathrm{loc}}>0$ is shown by the blue outer +/+/-+ / - symbols on the left and the labeled waveform on the right, producing an initial state close to the metastable state of $H$. The field for a small negative value $\beta\propto\Delta_{\mathrm{loc}}<0$ is shown by the teal inner +/+/-+ / - symbols on the left and the labeled waveform on the right, producing an initial state close to the ground state of $H$.
We study the long-lived oscillations of metastable and ground states in 1D neutral atom systems, using the setup shown in Figure 2. The atoms lie in a ring with $n_{s}$ sites, where we choose $n_{s}=16$. In particular, we use a staggered local detuning field to achieve confinement, yielding a confinement energy density $\mathcal{E}\propto\Delta_{\mathrm{loc}}$. We set the initial state $\left\lvert\psi(0)\right\rangle$ to be the ground state of $H_{\mathrm{prep}}=H(\beta_{\mathrm{prep}})$, obtained from exact diagonalization, where $H$ is the Rydberg Hamiltonian in Equation 1 and $\beta_{\mathrm{prep}}=-10^{-3}$. It is dominated by the $Z_{2}$ product state $\left\lvert 10\ldots 10\right\rangle$, represented by the colors in the atom chain in Figure 2. We evolve the state under $H$ with a different local detuning parameter $\beta$, which has a small constant amplitude $|\beta|\ll 1$. In effect, we perform a quench $\delta\beta=\beta-\beta_{\mathrm{prep}}$. The initial state is then close to the metastable state of $H$ for $\beta\propto\Delta_{\mathrm{loc}}>0$, shown by the blue outer +/+/-+ / - pattern and waveform in Figure 2; it is close to the ground state of $H$ for $\beta\propto\Delta_{\mathrm{loc}}<0$, shown by the teal inner +/+/-+ / - pattern and waveform. We note that in our procedure here, we hold the initial state fixed and modify the spatial pattern to obtain proximity to the metastable and ground states; this is the reverse of the schematic in Figure 1, where the spatial pattern is held fixed and the different states show the proximity.
@@
Figure 3: The oscillation properties of the magnetization $M$ for $\beta=\pm 0.01$. (a) The magnetization $M$ vs (dimensionless) time $\Omega t/2\pi$ for $\beta=0.01$. (b) The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ of the transformed magnetization $M_{T}$, defined in Equation 8, for $\beta=0.01$, which yields an initial state close to the metastable state. The curves with different opacities show results for $M_{T}$ evolved to different stop times $t_{\mathrm{Stop}}$. (c) The simultaneous eigenstates of $H$ and $T_{2}$. The eigenstates are identical for $\beta=\pm 0.01$, by symmetry. The ground state energy is set to zero, $E_{g}\rightarrow 0$, as indicated in the bottom left. The color bar shows the domain wall density $\rho n_{s}$; the blue points show the metastable state (MS) and ground state (GS), and the red points show the $1$-meson eigenstates. (d) The Fourier magnitude $|\tilde{M}_{T}(\omega)|$ for $\beta=-0.01$, which yields an initial state close to the ground state. In (b) and (d), the vertical grey lines show the frequencies $\omega_{\ell v}=E_{\ell}-E_{v}$, where $E_{\ell}$ for $\ell=1,\ldots,8$ are the energy eigenvalues of the 1-meson eigenstates with momentum $k=0$ and $E_{v}$ is the energy eigenvalue of the metastable or ground state $v=m,g$. The Fourier spectra for $\beta=\pm 0.01$ both have a similar dominant peak but different clusters of secondary peaks, which can be used to distinguish between the metastable and ground states, and understood from our analysis in Section V.